The mass of a piece of cork is 20g. What fraction of the cork is immersed when it floats in water?
(Take density of cork to be 0.25g/cm³)
1
Expert's answer
2018-12-17T08:01:10-0500

F_grav=F_Arch

F_grav=ρVg

F_Arch=ρ_water V_immersed·g

ρVg=ρ_water V_immersed·g

V_immersed=ρV/ρ_water =0.25V/1=1/4 V